OpenStudy (anonymous):

find three consecutive odd integers such that the sum of all three is 42 less than the product of the larger two. NEED HELP???

Directrix (directrix):

Let x be an odd integer The, the next consecutive odd integer is x + 2 And, the next is x + 4. Convert the words from the problem into symbols.

Directrix (directrix):

the sum of all three is 42 less than the product of the larger two. x + x + 2 + x + 4 = (x + 2)*(x + 4) -42 x + x + 2 + x + 4 = (x + 2)*(x + 4) - 42 @waffles56 The first task is to solve the above equation for x.

OpenStudy (anonymous):

THANK YOU @Directrix

Directrix (directrix):

You are welcome. Did you solve for the numbers?

OpenStudy (anonymous):

yes i got 5,7,9